OpenStudy (anonymous):

A substance has a half-life of 46 years. What percent of the original amount of the substance will remain after 20 years? Round to the nearest percent.

OpenStudy (anonymous):

44%

OpenStudy (anonymous):

Can you please show me step by step how to do that?

OpenStudy (anonymous):

Well, actually I am not quite sure but following my instincts I would think to find the percent, you would need to divide 20 by 46 20/46 = 0.43478260869... I would round that so 0.44 So that's 44% So if 44% is used up then 56% of the original substance is still available Not sure if this is the right way to do it but it seems logical :)

OpenStudy (anonymous):

I know I use the equation Po/2 = (Po)e^kt. I found a similar problem that says: "A substance has a half-life of 56 years. What percent of the original amount of the substance will remain after 20 years? Round to the nearest percent." And the answer was 78%. If I just do 20/56 I get 0.357, which would not be 78%. I sense I'm not understanding the equation itself. On the test I put the above equation, which the teacher said was right. But when I do... 46/2 = 46e^k20, I was wrong.

OpenStudy (anonymous):

This is all very confusing to me. What grade are you in? I'm in 9th

OpenStudy (anonymous):

This is college pre-calculus.

OpenStudy (anonymous):

Well that might explain why I don't understand it!

OpenStudy (anonymous):

it is not right i don't think

OpenStudy (anonymous):

try computing \[\left(\frac{1}{2}\right)^{\frac{20}{46}}\]

OpenStudy (anonymous):

maybe you get \(.44\) i don't know but it seems unlikely

OpenStudy (anonymous):

=(1/2)^20/46 =0.5^0.43 =0.74 Let me try with the other sample problem to see if I get the right answer.

OpenStudy (anonymous):

(1/2)^20/56 =0.5^ 0.36 =0.78= 78% Yeah, looks right. Now I need to think of it as an actual equation...

OpenStudy (anonymous):

i used only the number given

OpenStudy (anonymous):

half life is \(46\) so i used \(\left(\frac{1}{2}\right)^{\frac{t}{46}}\) for \(t=20\)

OpenStudy (anonymous):

I think I'm having trouble imagining it as Po/2=Poe^kt, which I'm not going to get credit for unless I can relate it to that. Hmm...

OpenStudy (anonymous):

we can do that too, it takes tons more work, but is not that hard

OpenStudy (anonymous):

half life is \(46\) years which means you can set \[e^{k\times 46}=\frac{1}{2}\] to solve for \(k\)

OpenStudy (anonymous):

Okay, so Po is just treated as "1" since it's the initial year, right?

OpenStudy (anonymous):

in equivalent logarithmic form you get \[46x=\ln(\frac{1}{2})\] and so \[k=\frac{\ln(.5)}{46}\]

OpenStudy (anonymous):

the \(P_0\) makes no difference if you start with 10 after 46 years you have 5 if you start with 1000 after 46 years you have 500 if you start with \(P_0\) after 46 years you have \(\frac{1}{2}P_0\) so first step in solving \[P_0e^{46k}=\frac{1}{2}P_0\] is to divide by \(P_0\) and start at \[e^{46k}=\frac{1}{2}\]

OpenStudy (anonymous):

notice it asks for "what percent remains" and doesn't tell you what you start with in any case the initial amount is unimportant